OpenStudy (anonymous):

can someone help me w/ a fraction problem.. 2(1/6)+d=1(1/9)

OpenStudy (studygurl14):

\[2(\frac{ 1 }{ 6 }) + d = 1(\frac{ 1 }{ 9 })\] First step. What is 1 times a number equal? The number. So \[1(\frac{ 1 }{ 9 }) = \frac{ 1 }{ 9 }\] Right? What is 2(1/6)? All you do is times the 2 to the 1 in the numerator, and get \[\frac{ 2 }{ 6 }\] What next?